# IDEA: Image Description Enhanced CLIP-Adapter

### v1

@@ Abstract
CLIP (Contrastive Language-Image Pre-training) has attained great success in pattern recognition and computer vision. Transferring CLIP to downstream tasks (e.g. zero- or few-shot classification) is a hot topic in multimodal learning. However, current studies primarily focus on either prompt learning for text or adapter tuning for vision, without fully exploiting the complementary information and correlations among image-text pairs. In this paper, we propose an Image Description Enhanced CLIP-Adapter (IDEA) method to adapt CLIP to few-shot image classification tasks. This method captures fine-grained features by leveraging both visual features and textual descriptions of images. IDEA is a training-free method for CLIP, and it can be comparable to or even exceeds state-of-the-art models on multiple tasks. Furthermore, we introduce Trainable-IDEA (T-IDEA), which extends IDEA by adding two lightweight learnable components (i.e., a projector and a learnable latent space), further enhancing the model’s performance and achieving SOTA results on 11 datasets. As one important contribution, we employ the Llama model and design a comprehensive pipeline to generate textual descriptions for images of 11 datasets, resulting in a total of 1,637,795 image-text pairs, named “IMD-11.” Our code and data are released at https://github.com/FourierAI/IDEA.
@@ 1 Introduction
Figure 1: Comparison of different PEFTs for CLIP. CoOp [12] and CoCoOp [13] have a similar architecture. Tip-Adapter [15], an extended version, shares the same architecture as CLIP-Adapter [14]. Different from previous works, (T)-IDEA introduces a multimodal adapter that explores the complementary relationship and semantic correlation among image-text pairs.
@@
    We design a comprehensive pipeline to generate image descriptions for 11 public image datasets, resulting in a total of 1,637,795 image-text pairs. Our dataset, which is referred to as “IMD-11”, has been made publicly available.
@@ 3.2 Image Description Enhanced CLIP-Adapter
where $\alpha\in[0,1]$ is a hyperparameter to balance the similarity between visual modal and textual modal, and $\beta\in(0,\infty)$ is a hyperparameter to trade off the few-shot knowledge and the zero-shot knowledge. The activation function $f(x)=\text{exp}(\theta(x-1))$ is defined for mapping the value of similarity to the interval [0,1]. $\theta\in(0,\infty)$ controls the sharpness of the activation function, which dynamically stretches and compresses the value of similarity to better fuse the few-shot knowledge to the zero-shot knowledge. Given the multimodal similarity among samples $\mathbf{X}\in\mathbb{R}^{NK\times 1}$, we define a function $g(\mathbf{X})=\sum_{K}\text{reshape}(\mathbf{X},N,K)$ to aggregate the similarity among samples to form the few-shot knowledge. $g(\mathbf{X})$ reshapes $\mathbf{X}$ into a matrix with $N$ rows and $K$ columns. We then sum the matrix in the column dimension. This operation is used to aggregate the instance-level similarity into the class-level similarity. Algorithm 1 shows the process of IDEA.
Algorithm 1 Pytorch style pseudocode for IDEA
{minted}
[linenos, fontfamily=tt,fontsize=, baselinestretch=1,tabsize=4]python # test image i_test D*1 # images in training set I_train NK*D # text in training set T_train NK*D # textual features for each class T_class N*D # hyperparameters alpha, beta, theta
# compute vision similarity sim_I = I_train @ i_test # compute text similarity sim_T = T_train @ i_test # compute multimodal similarity NK*1 sim = (1-alpha) * sim_I + alpha * sim_T # compute the activation function f(x) for multimodal similarity sim = torch.exp(theta(sim-1)) # aggregate similarity to form few-shot knowledge by g(X) few_shot_knowledge = torch.sum(sim.reshape(N,K),dim=1) # compute zero-shot knowledge N*1 zero_shot_knowledge = T_class @ i_test # compute logits N*1 logits = beta*few_shot_knowledge + zero_shot_knowledge

### v2

@@ Abstract
CLIP (Contrastive Language-Image Pre-training) has attained great success in pattern recognition and computer vision. Transferring CLIP to downstream tasks (e.g. zero- or few-shot classification) is a hot topic in multimodal learning. However, current studies primarily focus on either prompt learning for text or adapter tuning for vision, without fully exploiting the complementary information and correlations among image-text pairs. In this paper, we propose an Image Description Enhanced CLIP-Adapter (IDEA) method to adapt CLIP to few-shot image classification tasks. This method captures fine-grained features by leveraging both visual features and textual descriptions of images. IDEA is a training-free method for CLIP, and it can be comparable to or even exceeds state-of-the-art models on multiple tasks. Furthermore, we introduce Trainable-IDEA (T-IDEA), which extends IDEA by adding two lightweight learnable components (i.e., a projector and a learnable latent space), further enhancing the model’s performance and achieving SOTA results on 11 datasets. As one important contribution, we employ the Llama model and design a comprehensive pipeline to generate textual descriptions for images of 11 datasets, resulting in a total of 1,637,795 image-text pairs, named ”IMD-11”. Our code and data are released at https://github.com/FourierAI/IDEA.
@@
{highlights}
A new training-free adapter termed as IDEA is proposed for CLIP.
As a first attempt, IDEA explores complement and correlation information among image-text pairs.
IDEA achieves state-of-the-art performance in extensive experiments.
A new multimodal dataset is released to the community.
@@ 1 Introduction
Figure 1: Comparison of different PEFTs for CLIP. CoOp [12] and CoCoOp [13] have a similar architecture. Tip-Adapter [15] shares the same architecture as CLIP-Adapter [14]. Different from previous works, (T)-IDEA introduces a multimodal adapter that explores the complementary relationship and semantic correlation among image-text pairs.
@@
    We design a comprehensive pipeline to generate image descriptions for 11 public image datasets, resulting in a total of 1,637,795 image-text pairs. Our dataset, which is referred to as ”IMD-11”, has been made publicly available.
@@ 3.2 Image Description Enhanced CLIP-Adapter
where $\alpha\in[0,1]$ is a hyperparameter to balance the similarity between visual modal and textual modal, and $\beta\in(0,\infty)$ is a hyperparameter to trade off the few-shot knowledge and the zero-shot knowledge. The activation function $f(x)=\text{exp}(\theta(x-1))$ is defined for mapping the value of similarity to the interval [0,1]. $\theta\in(0,\infty)$ controls the sharpness of the activation function, which dynamically stretches and compresses the value of similarity to better fuse the few-shot knowledge to the zero-shot knowledge. Given the multimodal similarity among samples $\mathbf{X}\in\mathbb{R}^{NK\times 1}$, we define a function $g(\mathbf{X})=\sum_{K}\text{reshape}(\mathbf{X},N,K)$ to aggregate the similarity among samples to form the few-shot knowledge. $g(\mathbf{X})$ reshapes $\mathbf{X}$ into a matrix with $N$ rows and $K$ columns. We then sum the matrix in the column dimension. This operation is used to aggregate the instance-level similarity into the class-level similarity. Fig. 3 shows pseudocode of the core of an implementation of IDEA.
Figure 3: Pytorch style pseudocode for IDEA